OpenStudy (anonymous):

On Friday and Sat. There Were a total of 200 cars in the parking lot of a movie theater. On friday 120 cars were in the parking lot. a.what percents of total number of cars were in the parking lot on friday? b.what percent of total number of cars were in the parking lot on Saturday?

OpenStudy (anonymous):

a. = \[\frac{120}{200} \times 100% \]

OpenStudy (anonymous):

\[\frac{120}{200} \times 100\]

OpenStudy (anonymous):

b = \[\frac{80}{200} \times 100\]

OpenStudy (anonymous):

Thanks